In a story on National Public Radios show Morning Report (Airlines Use New Tech to Fight High Fuel Costs, Monday, March 14At that time Southwest Airlines operated about 200 B-737 airplanes. If their MARR was 12%, and assuming all of the modificati
Answer #1

(I) Let breakeven price be C per unit.

100,000 x C = $700,000 / P/A(12%, 10)

C = $7 / 5.6502

C = $1.24

(II) If C = $3.50,

PV ($) = - 700,000 + 100,000 x 3.5 x P/A(12%, 10)

= - 700,000 + 350,000 x 5.6502

= - 700,000 + 1,977,570

= 1,277,570